OpenStudy (anonymous):

(i square root of 10 - 4) (i square root of 10 + 4) i as in imaginary i not variable

OpenStudy (anonymous):

\[\left( i \sqrt{10}-4 \right)\left( i \sqrt{10} +4\right)\]

OpenStudy (phi):

I would use (a+b)(a-b)= a^2 - b^2

OpenStudy (anonymous):

would you mind explaining a bit more?

OpenStudy (phi):

do you know how to multiply two binomials (some people use FOIL) ?

OpenStudy (anonymous):

oh yes sorry I was confused for a second

OpenStudy (phi):

using just (a+b)(a-b) we get a^2 -ab +ab-b^2 which simplifies to a^2 - b^2 (we usually remember this as how to factor a *difference of squares* ) you can use FOIL, or you can use the "short-cut" \[ \left( i \sqrt{10}-4 \right)\left( i \sqrt{10} +4\right) = ( i \sqrt{10})^2 - 4^2 \]

OpenStudy (phi):

now you need to know what i*i is and what \( \sqrt{10} \cdot \sqrt{10} \) is

OpenStudy (anonymous):

i*i is -1 \[\sqrt{10} * \sqrt{10} is 10\]

OpenStudy (phi):

ok, now simplify \[ ( i \sqrt{10})^2 - 4^2 \]

OpenStudy (anonymous):

-26

OpenStudy (phi):

yes

OpenStudy (anonymous):

thank you